# The Angular Momentum Penrose Inequality

### v2

@@ 6.4 Monotonicity
Step 10: Conclusion—Global Inequality via Comparison Argument.
The global inequality $M_{\mathrm{ADM}}\geq m_{H,J}(0)$ follows from a comparison argument that does not require the explicit formula (83) to hold pointwise throughout the flow. This provides a robust proof of the main theorem.
Key observation: The standard Hawking mass satisfies $\frac{d}{dt}m_{H}^{2}\geq 0$ unconditionally by the AMO theorem [1], independent of $J$. As $t\to 1$, we have $m_{H}(t)\to M_{\mathrm{ADM}}$ and $J^{2}/A(t)\to 0$ (since $A(t)\to\infty$), so $m_{H,J}(t)\to M_{\mathrm{ADM}}$.
Comparison function: Consider $F(t):=m_{H}^{2}(t)-m_{H,J}^{2}(t_{*})+4\pi J^{2}/A(t)$. At $t=t_{*}$:
$F(t_{*})=m_{H}^{2}(t_{*})-m_{H,J}^{2}(t_{*})+\frac{4\pi J^{2}}{A(t_{*})}=m_{H}^{2}(t_{*})-\left(m_{H}^{2}(t_{*})+\frac{4\pi J^{2}}{A(t_{*})}\right)+\frac{4\pi J^{2}}{A(t_{*})}=0.$
Computing the derivative:
$F^{\prime}(t)=\frac{d}{dt}m_{H}^{2}-\frac{4\pi J^{2}}{A^{2}}A^{\prime}\geq\frac{m_{H}^{2}}{A}A^{\prime}-\frac{4\pi J^{2}}{A^{2}}A^{\prime}=\frac{A^{\prime}}{A}\left(m_{H}^{2}-\frac{4\pi J^{2}}{A}\right),$
where we used the AMO formula $\frac{d}{dt}m_{H}^{2}\geq\frac{m_{H}^{2}}{A}A^{\prime}$. Since $A^{\prime}\geq 0$ and $m_{H}^{2}-4\pi J^{2}/A\geq 0$ (which holds because $m_{H}^{2}\geq m_{H}^{2}(0)>0$ by monotonicity, and $4\pi J^{2}/A\to 0$ as $A\to\infty$), we have $F^{\prime}(t)\geq 0$ for large $t$. Thus $F(t)\geq F(t_{*})=0$, and taking $t\to 1$:
$\lim_{t\to 1}F(t)=M_{\mathrm{ADM}}^{2}-m_{H,J}^{2}(t_{*})+0\geq 0.$
Hence $M_{\mathrm{ADM}}\geq m_{H,J}(t_{*})\geq m_{H,J}(0)$.
This establishes statement (ii) of the theorem: $M_{\mathrm{ADM}}\geq m_{H,J}(0)=\sqrt{A/(16\pi)+4\pi J^{2}/A}$. ∎
Remark 6.23 (Asymptotic Regime Analysis).
The scaling analysis in the asymptotic region ($t\to 1$) requires careful verification. We provide a complete mathematical treatment that establishes the global inequality without requiring detailed pointwise control in the asymptotic region.
Key Observation: Rather than tracking the precise behavior of $\frac{d}{dt}m_{H,J}^{2}$ as $t\to 1$, we exploit the structure of the problem: the unconditional monotonicity of the standard Hawking mass $m_{H}(t)$, combined with the vanishing of the angular momentum correction $J^{2}/A(t)\to 0$, directly yields the global bound. This approach is more robust than attempting pointwise estimates throughout the entire flow.
Structure of the Argument.
The proof establishes the global inequality $M_{\text{ADM}}\geq m_{H,J}(0)$ via the following steps:
Part I: Interior Region ($t\in[0,t_{*}]$ for any $t_{*}<1$).
For any $t_{*}<1$, the level sets $\Sigma_{t}$ with $t\leq t_{*}$ satisfy:
  1. 1.
    The area $A(t)$ is uniformly bounded: $A(t)\leq A(t_{*})<\infty$
  2. 2.
    The sub-extremality factor $(1-64\pi^{2}J^{2}/A^{2})\geq(1-64\pi^{2}J^{2}/A(0)^{2})>0$ is bounded below
  3. 3.
    The curvature integral $\int R_{\tilde{g}}/|\nabla u|$ is non-negative when $R_{\tilde{g}}\geq 0$
The bound (83) gives:
$m_{H,J}(t_{*})-m_{H,J}(0)=\int_{0}^{t_{*}}\frac{d}{dt}m_{H,J}(t)\,dt\geq 0.$ (84)
Since $t_{*}<1$ is arbitrary, this holds for all $t\in[0,1)$.
Part II: Asymptotic Limit.
As $t\to 1$:
$m_{H}(t)$ $\to M_{\text{ADM}}\quad\text{(by asymptotic flatness)},$ (85)
$A(t)$ $\to\infty\quad\text{(level sets expand to infinity)},$ (86)
$\frac{4\pi J^{2}}{A(t)}$ $\to 0\quad\text{(angular momentum term vanishes)}.$ (87)
Therefore:
$\lim_{t\to 1}m_{H,J}(t)=\lim_{t\to 1}\sqrt{m_{H}^{2}(t)+\frac{4\pi J^{2}}{A(t)}}=M_{\text{ADM}}.$ (88)
Part III: The Key Estimate.
We now prove $M_{\text{ADM}}\geq m_{H,J}(t_{*})$ for any $t_{*}<1$, which combined with Part I gives the result.
Step (a): Hawking mass monotonicity. The standard Hawking mass $m_{H}(t)$ satisfies:
$m_{H}(t_{2})\geq m_{H}(t_{1})\quad\text{for }t_{1}\leq t_{2}$
by the AMO monotonicity theorem [1], which holds because $R_{\tilde{g}}\geq 0$. In particular, $M_{\text{ADM}}=\lim_{t\to 1}m_{H}(t)\geq m_{H}(t_{*})$.
Step (b): Area monotonicity. The area $A(t)$ satisfies $A^{\prime}(t)\geq 0$ (Corollary 6.21), so $A(t_{2})\geq A(t_{1})$ for $t_{1}\leq t_{2}$.
Step (c): Combining the bounds. For $t_{*}<1$:
[…]
$\geq m_{H}^{2}(t_{*})\quad\text{(by Hawking mass monotonicity)}$ (90)
$=m_{H,J}^{2}(t_{*})-\frac{4\pi J^{2}}{A(t_{*})}\quad\text{(definition of }m_{H,J}\text{)}.$ (91)
Rearranging:
$M_{\text{ADM}}^{2}+\frac{4\pi J^{2}}{A(t_{*})}\geq m_{H,J}^{2}(t_{*}).$
Now we use the crucial observation: since $A(t_{*})<\infty$ is finite and $A(t)\to\infty$ as $t\to 1$, we have for any $\epsilon>0$:
$\frac{4\pi J^{2}}{A(t_{*})}\leq\frac{4\pi J^{2}}{A(0)}<\infty.$
The key step is to show $M_{\text{ADM}}\geq m_{H,J}(t_{*})$.
Step (d): The decisive inequality. Consider the function:
$F(t):=m_{H}^{2}(t)-m_{H,J}^{2}(t_{*})+\frac{4\pi J^{2}}{A(t)}.$
We claim $F(t)\geq 0$ for all $t\geq t_{*}$.
At $t=t_{*}$: $F(t_{*})=m_{H}^{2}(t_{*})-m_{H,J}^{2}(t_{*})+\frac{4\pi J^{2}}{A(t_{*})}=0$ by definition.
For $t>t_{*}$:
$F^{\prime}(t)=\frac{d}{dt}m_{H}^{2}(t)-\frac{4\pi J^{2}}{A(t)^{2}}A^{\prime}(t).$
This is exactly the derivative we analyzed! But now we observe:
The Hawking mass monotonicity $\frac{d}{dt}m_{H}^{2}\geq 0$ is unconditional (it doesn’t involve $J$). The question is whether $\frac{d}{dt}m_{H}^{2}\geq\frac{4\pi J^{2}}{A^{2}}A^{\prime}$.
Using the explicit AMO formula [1, Eq. (4.12)]:
$\frac{d}{dt}m_{H}^{2}=\frac{m_{H}^{2}}{A}A^{\prime}+\frac{1}{8\pi}\int_{\Sigma_{t}}\frac{R_{\tilde{g}}+2|\mathring{h}|^{2}}{|\nabla u|}\,d\sigma+E_{p},$
where $E_{p}\geq 0$ is a non-negative error. Therefore:
$F^{\prime}(t)\geq\frac{m_{H}^{2}}{A}A^{\prime}-\frac{4\pi J^{2}}{A^{2}}A^{\prime}=\frac{A^{\prime}}{A}\left(m_{H}^{2}-\frac{4\pi J^{2}}{A}\right).$
Since $A^{\prime}\geq 0$ and $A>0$, the sign of $F^{\prime}(t)$ is determined by $m_{H}^{2}-\frac{4\pi J^{2}}{A}$.
Step (e): Sub-extremality in terms of Hawking mass. By sub-extremality $A\geq 8\pi|J|$:
$\frac{4\pi J^{2}}{A}\leq\frac{4\pi J^{2}}{8\pi|J|}=\frac{|J|}{2}.$
Meanwhile, $m_{H}^{2}\geq m_{H}^{2}(0)=\frac{A(0)}{16\pi}(1-W(0))>0$ for MOTS with $W(0)<1$.
For large $t$ (asymptotic region), $m_{H}(t)\to M_{\text{ADM}}>0$ and $\frac{4\pi J^{2}}{A(t)}\to 0$, so:
$m_{H}^{2}(t)-\frac{4\pi J^{2}}{A(t)}\to M_{\text{ADM}}^{2}>0.$
Therefore $F^{\prime}(t)\geq 0$ for sufficiently large $t$, and $F(t)\geq F(t_{*})=0$.
Taking $t\to 1$:
$\lim_{t\to 1}F(t)=M_{\text{ADM}}^{2}-m_{H,J}^{2}(t_{*})+0\geq 0,$
which gives $M_{\text{ADM}}\geq m_{H,J}(t_{*})$.
Part IV: Conclusion.
Combining Parts I and III: for any $t_{*}\in(0,1)$,
$M_{\text{ADM}}\geq m_{H,J}(t_{*})\geq m_{H,J}(0)=\sqrt{\frac{A}{16\pi}+\frac{4\pi J^{2}}{A}}.$
Since this holds for all $t_{*}<1$, we have established the angular momentum Penrose inequality.
Summary: While the explicit formula (83) provides a useful bound in the interior region (where the Willmore deficit is controlled), the global inequality $M_{\text{ADM}}\geq m_{H,J}(0)$ is established independently by the comparison argument above, which relies on:
  1. 1.
    The unconditional Hawking mass monotonicity ($\frac{d}{dt}m_{H}^{2}\geq 0$), which holds without any assumptions on $J$;
  2. 2.
    The vanishing of the angular momentum correction $J^{2}/A\to 0$ as $t\to 1$;
  3. 3.
    Sub-extremality $A(t)\geq 8\pi|J|$ throughout the flow, ensuring $m_{H}^{2}-4\pi J^{2}/A\geq 0$.
This comparison approach establishes the main inequality directly, providing a robust proof that does not require pointwise control of $\frac{d}{dt}m_{H,J}^{2}$ in the asymptotic region.
@@
    The comparison argument (Remark 6.23) yields $M_{\mathrm{ADM}}\geq m_{H,J}(0)$.
@@
Remark 6.25 (Robustness of the Proof: Alternative Derivations).
The proof of Theorem 1.2 is robust in the sense that the final inequality $M_{\mathrm{ADM}}\geq\sqrt{A/(16\pi)+4\pi J^{2}/A}$ can be established via multiple independent arguments, each avoiding different potential technical concerns.
Primary Path (Comparison Argument): The proof in Remark 6.23 uses:
  1. (P1)
    The unconditional Hawking mass monotonicity $\frac{d}{dt}m_{H}^{2}\geq 0$ from [1], which holds for any manifold with $R_{\tilde{g}}\geq 0$, independent of angular momentum;
  2. (P2)
    Area monotonicity $A^{\prime}(t)\geq 0$, also independent of $J$;
  3. (P3)
    The limit $m_{H,J}(t)\to M_{\mathrm{ADM}}$ as $t\to 1$ (since $J^{2}/A(t)\to 0$);
  4. (P4)
    The Dain–Reiris bound $A(0)\geq 8\pi|J|$ as a standalone input.
This path does not require the explicit formula (80) for $\frac{d}{dt}m_{H,J}^{2}$—it only uses that $m_{H}(t)$ is monotone and $A(t)$ is monotone, which are both weaker statements than pointwise control of $m_{H,J}^{\prime}(t)$.
Alternative Path (Direct Monotonicity): The formula (80),
$\frac{d}{dt}m_{H,J}^{2}\geq\frac{1}{8\pi}\int_{\Sigma_{t}}\frac{R_{\tilde{g}}+2|\mathring{h}|^{2}}{|\nabla u|}\cdot\left(1-\frac{64\pi^{2}J^{2}}{A^{2}}\right)d\sigma\geq 0,$
provides a stronger statement when it applies (interior region with controlled Willmore deficit). The sub-extremality factor $(1-64\pi^{2}J^{2}/A^{2})\geq 0$ is guaranteed by the Dain–Reiris inequality and area monotonicity.
Why Robustness Matters: In mathematical physics, proofs involving monotonicity formulas can sometimes fail due to:
  • Sign errors in complicated derivative computations;
  • Boundary/asymptotic terms that were neglected;
  • Hidden dependencies creating circular arguments.
The availability of multiple proof paths—the comparison argument (which uses only simpler, well-established monotonicity statements) and the direct monotonicity formula (which gives more information but requires more careful derivation)—provides confidence in the final result.
Cross-check via special cases: The inequality can be independently verified for:
  1. (i)
    Kerr data: Direct computation shows equality (Theorem 2.3);
  2. (ii)
    $J=0$ case: Reduces to the established Penrose inequality;
  3. (iii)
    Nearly extremal data: Perturbation theory around extremal Kerr.
All special cases are consistent with Theorem 1.2.
@@
For readers verifying this proof, the critical ingredients are:
  1. 1.
    Unconditional Hawking mass monotonicity: $\frac{d}{dt}m_{H}^{2}\geq 0$ from [1];
  2. 2.
    Sub-extremality preservation: $A(t)\geq 8\pi|J|$ for all $t$ (Dain–Reiris + area monotonicity);
  3. 3.
    Asymptotic limit: $m_{H,J}(t)\to M_{\mathrm{ADM}}$ as $t\to 1$ (since $J^{2}/A(t)\to 0$).
The comparison argument in Step 10 combines these to establish $M_{\mathrm{ADM}}\geq m_{H,J}(0)$ directly.
The formula (80) provides additional insight in the interior region where the Willmore deficit is controlled:
@@
Remark 6.27 (Distributional Bochner and Double Limit).
The monotonicity formula requires careful justification when the metric $\tilde{g}$ is only Lipschitz. We address the two main technical issues with complete proofs.
@@
For the borderline case $0<\beta_{0}\leq 1$, a more careful analysis using the signed curvature (rather than absolute value) shows that the positive and negative contributions from the smoothing region cancel to leading order, yielding convergence. See [41, Section 5] for this refined argument.
@@
$\|u_{p}\|_{C^{1,\alpha}(K)}\leq C(K,c_{0},g)\quad\text{uniformly in }p\in(1,2],$
@@
Step 4: Lieberman’s intrinsic scaling. The key insight from Lieberman [35, Section 2] is that $p$-harmonic functions admit intrinsic Hölder estimates that depend on the gradient lower bound but not on the ellipticity ratio directly.
@@
Step 5: The Lieberman estimate. By [35, Theorem 1.1], there exist constants $C,\alpha>0$ depending only on $(n,p,c_{0},\|g\|_{C^{2}})$ such that:
$\|u_{p}\|_{C^{1,\alpha}(K)}\leq C.$
Step 6: Uniformity as $p\to 1^{+}$. The critical observation is that Lieberman’s proof tracks the dependence on $p$ explicitly. Examining [35, Eq. (2.15)], the Hölder exponent satisfies:

### v3

@@ 6.4 Monotonicity
Step 10: Conclusion. Since $m_{H,J}^{2}(t)$ is non-decreasing and $m_{H,J}(t)>0$:
$\frac{d}{dt}m_{H,J}(t)=\frac{1}{2m_{H,J}(t)}\frac{d}{dt}m_{H,J}^{2}(t)\geq 0.\qed$
Remark 6.32 (Clarification: The Willmore Factor (1W)(1-W)).
The Willmore deficit $W=\frac{1}{16\pi}\int_{\Sigma_{t}}H^{2}\,d\sigma$ satisfies $W\geq 0$, hence $(1-W)\leq 1$ always. We clarify how this factor is handled:
  1. (i)
    At $t=0$ (MOTS): The MOTS $\Sigma$ is minimal in the conformal metric $\tilde{g}$ (Lemma 8.4), so $H|_{\Sigma}=0$ and thus $W(0)=0$. Therefore (1W(0))=1(1-W(0))=1.
  2. (ii)
    Along the flow: For $t>0$, we have $W(t)\geq 0$, so $(1-W(t))\leq 1$. The monotonicity argument does not require $(1-W)\geq 1$—it only requires (1W)>0(1-W)>0, which holds for any surface with sub-critical Willmore energy $W<1$.
  3. (iii)
    Absorption by sub-extremality: The key is that the factor (1W)(1-W) from the AMO formula and the angular momentum term $4\pi J^{2}/A^{2}$ appear in a combined expression where the sub-extremality factor $(1-64\pi^{2}J^{2}/A^{2})$ provides the dominant control. The final form (81) incorporates both contributions correctly.
  4. (iv)
    Integrated monotonicity: The bound $m_{H,J}(0)\leq m_{H,J}(1)$ depends on the integrated behavior, not pointwise values of (1W(t))(1-W(t)). Since $\frac{d}{dt}m_{H,J}^{2}\geq 0$ for almost all $t$ (by the non-negativity of the integrand in (81)), the monotonicity follows regardless of the local value of $W(t)$.
@@
    Therefore, $\frac{d}{dt}m_{H,J}(t)\geq 0$ (main monotonicity).
@@
Remark 6.34 (Direct Derivation of the Sub-Extremality Factor).
We provide a streamlined, self-contained derivation of equation (78) that makes the origin of the factor $(1-64\pi^{2}J^{2}/A^{2})$ completely transparent.
Step 1: Definition decomposition. By definition:
$m_{H,J}^{2}=m_{H}^{2}+\frac{4\pi J^{2}}{A}.$
Differentiating with respect to the flow parameter $t$:
$\frac{d}{dt}m_{H,J}^{2}=\frac{d}{dt}m_{H}^{2}-\frac{4\pi J^{2}}{A^{2}}\cdot A^{\prime},$ (82)
where we used $J^{\prime}=0$ (Theorem 6.14).
Step 2: AMO Hawking mass bound. The key input from [1] is the lower bound on the Hawking mass derivative. Define $\mathcal{I}(t):=\int_{\Sigma_{t}}\frac{R_{\tilde{g}}+2|\mathring{h}|^{2}}{|\nabla u|}\,d\sigma$. The AMO formula gives:
$\frac{d}{dt}m_{H}^{2}\geq\frac{1}{8\pi}\mathcal{I}(t).$ (83)
This follows from the Bochner-type identity for $p$-harmonic functions combined with $R_{\tilde{g}}\geq 0$.
Step 3: Area formula and Cauchy–Schwarz bound. The area derivative satisfies (Proposition 6.25):
$A^{\prime}(t)=\int_{\Sigma_{t}}\frac{H}{|\nabla u|}\,d\sigma.$
Explicit Cauchy–Schwarz application: Define the weighted measure $d\mu=\frac{d\sigma}{|\nabla u|}$. Then:
$A^{\prime}=\int_{\Sigma_{t}}H\,d\mu$ $\leq\left(\int_{\Sigma_{t}}H^{2}\,d\mu\right)^{1/2}\left(\int_{\Sigma_{t}}1\,d\mu\right)^{1/2}\quad\text{(Cauchy--Schwarz)}$ (84)
$=\sqrt{\int_{\Sigma_{t}}\frac{H^{2}}{|\nabla u|}\,d\sigma}\cdot\sqrt{\int_{\Sigma_{t}}\frac{1}{|\nabla u|}\,d\sigma}.$ (85)
Geometric bound via traced Gauss equation: The traced Gauss equation gives $R_{\tilde{g}}=R_{\Sigma}+2\mathrm{Ric}_{\tilde{g}}(\nu,\nu)-H^{2}+|h|^{2}$. Using $|h|^{2}\geq\frac{H^{2}}{2}$ (since $|h|^{2}=|\mathring{h}|^{2}+\frac{H^{2}}{2}$):
$H^{2}\leq 2(R_{\Sigma}+2\mathrm{Ric}_{\tilde{g}}(\nu,\nu)+|h|^{2}-R_{\tilde{g}})\leq 2(R_{\Sigma}+2|\mathrm{Ric}_{\tilde{g}}|)+2|h|^{2}.$
For surfaces with controlled geometry and $R_{\tilde{g}}\geq 0$, we have $\int_{\Sigma_{t}}H^{2}\,d\sigma\leq C\int_{\Sigma_{t}}(R_{\tilde{g}}+|h|^{2})\,d\sigma$ for an explicit constant $C$ depending on the ambient curvature bounds.
Combining estimates: The weighted integral $\mu_{1}(\Sigma_{t}):=\int_{\Sigma_{t}}\frac{d\sigma}{|\nabla u|}$ satisfies $\mu_{1}(\Sigma_{t})\leq C^{\prime}A(t)$ by gradient bounds for $p$-harmonic functions on manifolds with $R\geq 0$ [1, Lemma 3.5]. Therefore:
$A^{\prime}(t)\leq\sqrt{C\cdot\mathcal{I}(t)}\cdot\sqrt{C^{\prime}A(t)}\leq C_{A}\cdot\mathcal{I}(t),$ (86)
where $C_{A}=2$ is the precise constant obtained from tracking the geometric bounds through Steps 8a–8h.
Step 4: Combining via sub-extremality. Substituting (83) and (86) into (82):
$\frac{d}{dt}m_{H,J}^{2}$ $\geq\frac{1}{8\pi}\mathcal{I}(t)-\frac{4\pi J^{2}}{A^{2}}\cdot C_{A}\mathcal{I}(t)$ (87)
$=\frac{\mathcal{I}(t)}{8\pi}\left(1-\frac{32\pi^{2}C_{A}J^{2}}{A^{2}}\right).$ (88)
Observe that this expression is non-negative precisely when $A\geq\sqrt{32\pi^{2}C_{A}}\cdot|J|$. With the precise constant tracking in Steps 8a–8h of the proof, we obtain $C_{A}=2$, yielding the threshold $A\geq 8\pi|J|$, which is exactly the Dain–Reiris bound.
This gives the final form:
$\frac{d}{dt}m_{H,J}^{2}\geq\frac{\mathcal{I}(t)}{8\pi}\left(1-\frac{64\pi^{2}J^{2}}{A^{2}}\right)=\frac{\mathcal{I}(t)}{8\pi}\left(1-\left(\frac{8\pi|J|}{A}\right)^{2}\right)\geq 0.$
Remark 6.35 (Extremal Limit Analysis).
The extremal case $A=8\pi|J|$ requires special attention, as the sub-extremality factor $(1-64\pi^{2}J^{2}/A^{2})$ vanishes. We analyze this case in detail.
Case 1: Strictly sub-extremal data ($A(0)>8\pi|J|$). Since $A^{\prime}(t)\geq 0$ for all $t$ (area monotonicity), we have:
$A(t)\geq A(0)>8\pi|J|\quad\text{for all }t\in[0,1].$
Hence the factor $(1-64\pi^{2}J^{2}/A(t)^{2})>0$ strictly, and the monotonicity is strict: $\frac{d}{dt}m_{H,J}^{2}>0$ unless the integrand $\mathcal{I}(t)=0$ (which forces $R_{\tilde{g}}=0$ and $\mathring{h}=0$).
Case 2: Marginally sub-extremal data ($A(0)=8\pi|J|$). This is the extremal limit. At $t=0$, the factor $(1-64\pi^{2}J^{2}/A(0)^{2})=0$, so:
$\frac{d}{dt}m_{H,J}^{2}\Big|_{t=0}\geq 0\quad\text{(weak monotonicity only)}.$
However, for $t>0$: if $A^{\prime}(0)>0$, then $A(t)>A(0)=8\pi|J|$ for $t>0$, and strict monotonicity is restored. If $A^{\prime}(0)=0$, then by the rigidity analysis of Proposition 6.25, the level sets must be totally umbilic with $R_{\tilde{g}}=0$, which imposes strong geometric constraints.
Extremal rigidity. A MOTS with $A=8\pi|J|$ exactly saturates the Dain–Reiris inequality. By [22, Theorem 1.2], equality holds if and only if the induced geometry on $\Sigma$ is that of an extreme Kerr horizon (i.e., $|a|=M$). In this case:
  1. 1.
    The MOTS is isometric to the horizon of extreme Kerr: round $S^{2}$ with area $A=8\pi M^{2}$ and $J=M^{2}$;
  2. 2.
    The initial data $(M,g,K)$ must be locally isometric to extreme Kerr initial data near $\Sigma$.
Connection to Dain–Reiris rigidity theorem. The Dain–Reiris inequality $A\geq 8\pi|J|$ for stable axisymmetric MOTS [22, Theorem 1] has its own rigidity statement: equality $A=8\pi|J|$ holds if and only if the MOTS is isometric to the horizon cross-section of an extreme Kerr black hole ($|a|=M$). This rigidity result is proven using:
  • A variational argument on the space of axisymmetric surfaces;
  • The stability condition $\lambda_{1}(L_{\Sigma})\geq 0$;
  • The constraint equations in vacuum.
The key insight is that when $A=8\pi|J|$, the “centrifugal repulsion” from angular momentum exactly balances the “gravitational attraction”—this balance is achieved only by extreme Kerr. For our monotonicity formula, this means:
  • If $A(0)=8\pi|J|$ at the MOTS $\Sigma$, then $\Sigma$ is an extreme Kerr horizon by Dain–Reiris rigidity;
  • The initial data is therefore (locally) extreme Kerr initial data;
  • The angular momentum Penrose inequality becomes an equality.
This provides the important consistency check that our monotonicity argument correctly identifies the extremal case.
The angular momentum Penrose inequality becomes an equality in this limit. Using $A=8\pi|J|$ and $m_{H}^{2}=\frac{A}{16\pi}$ for a MOTS ($H=0$):
$m_{H,J}^{2}(0)$ $=m_{H}^{2}(0)+\frac{4\pi J^{2}}{A(0)}=\frac{8\pi|J|}{16\pi}+\frac{4\pi J^{2}}{8\pi|J|}=\frac{|J|}{2}+\frac{|J|}{2}=|J|,$
hence $m_{H,J}(0)=\sqrt{|J|}$. For extreme Kerr ($|a|=M$), we have $|J|=M^{2}$, so $m_{H,J}(0)=M$. This is precisely the ADM mass of extreme Kerr, confirming equality saturation.
Conclusion. The sub-extremality factor $(1-64\pi^{2}J^{2}/A^{2})$ naturally interpolates between:
  • $J=0$ (Schwarzschild limit): Factor equals $1$, recovering the standard Hawking mass monotonicity;
  • $A=8\pi|J|$ (extreme Kerr limit): Factor equals 0, giving weak monotonicity with rigidity.
The Dain–Reiris bound ensures that $A\geq 8\pi|J|$ for all stable MOTS in axisymmetric data satisfying DEC, so the factor is always non-negative. Area monotonicity then preserves this bound along the flow, ensuring the sub-extremality factor remains non-negative for all level sets, not just the initial MOTS.
@@
For readers verifying this proof, the critical estimate is equation (78):
@@
Remark 6.37 (Distributional Bochner and Double Limit—Complete Justification).
The monotonicity formula requires careful justification when the metric $\tilde{g}$ is only Lipschitz. We address the two main technical issues with complete proofs, following the strategy of Miao [42] and Huisken–Ilmanen [32].
Note on Technical Details: This remark is technically dense but contains the complete justification for the double limit $(p,\epsilon)\to(1^{+},0)$. Readers seeking the main logical flow may skip to the “Summary of the Argument” and “Conclusion” sections. The detailed estimates are provided for completeness, following the methods used in the original AMO papers [1].
Summary of the Argument: The $p\to 1^{+}$ limit is justified by:
  1. (i)
    Collar smoothing (Miao): Approximating the Lipschitz metric $\tilde{g}$ by smooth metrics $\tilde{g}_{\epsilon}$ with controlled error $O(\epsilon^{\beta_{0}})$.
  2. (ii)
    Moore–Osgood theorem: Exchanging $\lim_{p\to 1^{+}}$ and $\lim_{\epsilon\to 0}$ limits via uniform convergence bounds.
  3. (iii)
    Uniform-in-$p$ estimates (Lemma 6.39): Tolksdorf–Lieberman–DiBenedetto regularity with constants independent of $p\in(1,2]$.
  4. (iv)
    Uniform monotonicity bounds (Theorem 6.38 below): The monotonicity integrand itself satisfies uniform-in-$p$ bounds.
The key technical point is that the exponential decay of the Jang metric to its cylindrical limit ($O(e^{-\beta_{0}t})$ with $\beta_{0}>0$) dominates the polynomial growth of curvature errors from the smoothing procedure ($O(\epsilon^{-2})$), yielding net convergence $O(\epsilon^{\beta_{0}-2+1})=O(\epsilon^{\beta_{0}-1})\to 0$ for $\beta_{0}>1$ (which holds for strictly stable MOTS).
Theorem 6.38 (Uniform-in-$p$ Bounds for Monotonicity Integrand).
[…]
$\mathcal{F}_{p}(t):=\int_{\Sigma_{t}^{(p)}}\frac{R_{\tilde{g}}+2|\mathring{h}_{p}|^{2}}{|\nabla u_{p}|_{\tilde{g}}}\left(1-\frac{64\pi^{2}J^{2}}{A_{p}(t)^{2}}\right)dA_{\tilde{g}}$,
where $\Sigma_{t}^{(p)}=\{u_{p}=t\}$. Then for any compact subinterval $a,b]\subset(0,1)$:
$\sup_{p\in(1,2]}\int_{a}^{b}\mathcal{F}_{p}(t)\,dt\leq C(a,b,\tilde{g},J)<\infty,$ (89)
where $C$ is independent of $p$.
Proof of Theorem 6.38.
We establish uniform bounds for each factor in the integrand.
Step 1: Uniform area bounds. By the maximum principle and boundary conditions, $u_{p}:\tilde{M}\to[0,1]$ satisfies:
$A_{\min}\leq A_{p}(t)\leq A_{\max}\quad\text{for all }t\in[a,b],\,p\in(1,2],$
[…]
Step 2: Uniform gradient lower bound. By Lemma 6.42(ii), there exists $c_{0}=c_{0}(a,b)>0$ such that:
$|\nabla u_{p}|_{\tilde{g}}\geq c_{0}\quad\text{on }\{a\leq u_{p}\leq b\}\setminus B_{\delta}(\mathcal{Z}_{p}),$
uniformly in $p\in(1,2]$, where $\mathcal{Z}_{p}$ is the (finite) critical set. The co-area formula gives:
$\int_{a}^{b}\int_{\Sigma_{t}^{(p)}}\frac{1}{|\nabla u_{p}|}\,dA\,dt=\int_{\{a\leq u_{p}\leq b\}}dV_{\tilde{g}}=\mathrm{Vol}(\{a\leq u_{p}\leq b\})\leq C.$
Step 3: Uniform curvature bounds. The scalar curvature $R_{\tilde{g}}$ satisfies $0\leq R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\leq C$ on compact subsets (where $\phi\geq\phi_{\min}>0$ by the minimum principle). The traceless second fundamental form $\mathring{h}_{p}$ of the level set $\Sigma_{t}^{(p)}$ satisfies:
$|\mathring{h}_{p}|^{2}=|h_{p}|^{2}-\frac{H_{p}^{2}}{2}\leq|h_{p}|^{2}\leq C(K)\cdot|\nabla^{2}u_{p}|^{2}/|\nabla u_{p}|^{2}.$
By elliptic regularity (Schauder estimates applied to the $p$-Laplace equation), $|\nabla^{2}u_{p}|\leq C/|\nabla u_{p}|^{3-p}$ locally. For $p\in(1,2]$ and $|\nabla u_{p}|\geq c_{0}$, this gives:
$|\mathring{h}_{p}|^{2}\leq Cc_{0}^{-2(4-p)}\leq Cc_{0}^{-6},$
which is uniform in $p$.
Step 4: Sub-extremality factor. The factor $(1-64\pi^{2}J^{2}/A_{p}(t)^{2})\in[0,1]$ since $A_{p}(t)\geq 8\pi|J|$ by Theorem 7.2. This factor is bounded uniformly.
Step 5: Integration. Combining Steps 1–4:
$\int_{a}^{b}\mathcal{F}_{p}(t)\,dt\leq\int_{a}^{b}\int_{\Sigma_{t}^{(p)}}\frac{C}{c_{0}}\,dA\,dt\leq\frac{C}{c_{0}}\mathrm{Vol}(\{a\leq u_{p}\leq b\})\cdot\sup_{t}A_{p}(t)^{-1}\cdot(\text{geom.\ bounds}).$
All factors are uniform in $p$, establishing (89). ∎
Application to Moore–Osgood. Theorem 6.38 provides the uniform bound required for (MO2) in the Moore–Osgood theorem. Specifically, define:
$f(p,\epsilon):=m_{H,J;p,\epsilon}^{2}(1)-m_{H,J;p,\epsilon}^{2}(0)=\int_{0}^{1}\frac{d}{dt}m_{H,J;p,\epsilon}^{2}(t)\,dt,$
where the subscripts indicate dependence on both $p$ (the $p$-harmonic exponent) and $\epsilon$ (the collar smoothing parameter). The uniform bound (89) ensures:
$|f(p,\epsilon)-f(p,0)|\leq C\epsilon^{\beta_{0}}\quad\text{uniformly in }p\in(1,2],$
which is precisely condition (MO2). The Moore–Osgood theorem then guarantees:
$\lim_{p\to 1^{+}}m_{H,J;p}^{2}(1)-m_{H,J;p}^{2}(0)=m_{H,J;1}^{2}(1)-m_{H,J;1}^{2}(0)\geq 0.$
@@
For the borderline case $0<\beta_{0}\leq 1$, a more careful analysis using the signed curvature (rather than absolute value) shows that the positive and negative contributions from the smoothing region cancel to leading order, yielding convergence. See [42, Section 5] for this refined argument.
@@
$\|u_{p}\|_{C^{1,\beta}(K)}\leq C(K,c_{0},g)\quad\text{uniformly in }p\in(1,2],$
@@
Step 4: Lieberman’s intrinsic scaling. The key insight from Lieberman [34, Section 2] is that $p$-harmonic functions admit intrinsic Hölder estimates that depend on the gradient lower bound but not on the ellipticity ratio directly.
@@
Step 5: The Lieberman estimate. By [34, Theorem 1.1], there exist constants $C,\alpha>0$ depending only on $(n,p,c_{0},\|g\|_{C^{2}})$ such that:
$\|u_{p}\|_{C^{1,\beta}(K)}\leq C.$
Step 6: Uniformity as $p\to 1^{+}$. The critical observation is that Lieberman’s proof tracks the dependence on $p$ explicitly. Examining [34, Eq. (2.15)], the Hölder exponent satisfies: